OpenStudy (anonymous):

Help. Simplify. x-3/5x-4 - 3(4-3x)/4-5x

OpenStudy (anonymous):

\[\Large \frac{x-3}{5x-4}-3\frac{4-3x}{4-5x}\] did i write it correctly? if its correct let me know.

OpenStudy (anonymous):

3(4-3x)

OpenStudy (anonymous):

ok lets proceed further \[\Large \frac{x-3}{5x-4}-\frac{3(4-3x)}{4-5x}\] we can write it as \[\Large \frac{x-3}{5x-4}+\frac{3(4-3x)}{5x-4}\] (just taking minus out of the denominator in the second fraction) now taking 5x-4 as LCM \[\Large \frac{x-3+3(4-3x)}{5x-4}\]

OpenStudy (anonymous):

can you further simplify it now?

OpenStudy (anonymous):

-8x+9/ 5x-4 right?

OpenStudy (anonymous):

yes thats right .